# On importance sampling and independent Metropolis–Hastings with an unbounded weight function

### v1

@@ 4.2 Polynomial convergence rates
Furthermore, if Assumption 2 also holds, there exists a constant $C$, independent of $t$ and $N$, such that for all $N\geq 1$ and $t\geq 0$,
@@
where the first inequality is Cauchy–Schwarz and the second uses Proposition 1.1 under Assumption 2. Furthermore, in the large $N$ asymptotics we expect $1-\hat{Z}(\mathbf{x})$ to behave as a Normal distribution with mean zero and standard deviation $\sqrt{q(\omega^{2})-1}/\sqrt{N}$, so that its expectation should indeed behave as $\sqrt{(2/\pi)(q(\omega^{2})-1)}/\sqrt{N}$.
@@
$\left|P^{t}(\mathbf{x},\cdot)-\bar{\pi}\right|_{\text{TV}}\leq(r(\mathbf{x}))^{t}+\frac{\tilde{C}}{\sqrt{N}(1+t)^{p-1}}.$ (30)
@@
$\left|P^{t}(x,\cdot)-\pi\right|_{\text{TV}}\leq(r(x))^{t}+\frac{D}{(1+t)^{p-1}}.$ (31)
The purpose of the following example is to demonstrate that the rate $t^{-(p-1)}$ in Corollary 4.1 and Proposition 4.4 cannot be improved beyond polylogarithmic factors. The proof is provided in Appendix A.4.4.
@@ 4.4 Comparing the biases of IS and IMH
We can now compare the bias of IMH and IS under the assumption $q(\omega^{p})<\infty$ for $p\geq 2$. Suppose that the goal is to obtain a single sample close to $\pi$ in total variation distance, with a budget of $N$ samples from $q$ and $N$ evaluations of $\omega$.
  • One approach is sampling-importance resampling (SIR), which refers to the following procedure. First, run Algorithm 1. Then, draw $k\sim\text{Categorical}(\omega(x_{1}),\ldots,\omega(x_{N}))$ and return $x_{k}$. For a test function $f$ with $|f|_{\infty}\leq 1$, under the conditions of Theorem 2.1 the marginal distribution $\mu_{N}^{\text{SIR}}$ of $x_{k}$ satifies
    $\mu_{N}^{\text{SIR}}(f)-\pi(f)\sim_{N\to\infty}-q(\omega^{2}\cdot(f-\pi(f)))N^{-1}.$ (34)
    Skare et al. (2003) provide a similar result in the case of bounded weights, and propose a modification of SIR to reduce this bias to $N^{-2}$.
  • On the other hand, Theorem 4.1 suggests that IMH (with one particle) after $N$ iterations provides a sample from a distribution $qP^{N-1}$, for which, under the condition $q(\omega^{p})<\infty$,
@@
    Thus, the bias may be reduced faster when increasing the computing budget in IMH rather than in SIR, as soon as $p>2$; for the modified SIR of Skare et al. (2003), IMH is faster as soon as $p>3$.
In the case of bounded weights, MCMC methods such as PIMH or particle Gibbs (Andrieu et al., 2010) are geometrically ergodic (e.g. Lee et al., 2020) and the bias comparison is clearly at the advantage of MCMC algorithms, as discussed in Cardoso et al. (2022).
@@ 5.1 Construction
The bias of importance sampling was described in Section 2, and that of IMH in Section 4. Here we consider the removal of the bias, and the associated cost. For this purpose, Middleton et al. (2019) employ common random numbers couplings of PIMH and the approach of Glynn & Rhee (2014). We pursue this strategy.
Upon running Algorithm 5 with $N\geq 1$, with $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t-1}\}$, one can compute the following unbiased estimator:
@@
(36) is an instance of unbiased MCMC (Jacob et al., 2020; Atchadé & Jacob, 2024), where here MCMC is PIMH, and various generic improvements could be considered, such as increasing the lag between the chains, or introducing a burn-in parameter. However, in the particular case of PIMH, the number of particles $N$ is a key parameter and here we focus on the regime $N\to\infty$, in which case $\hat{F}_{u}$ naturally compares with $\hat{F}(\mathbf{x}_{0})$, which is the regular IS estimator. Hence we call (36) the unbiased self-normalized importance sampling estimator (UIS), but it could also be called unbiased PIMH.
@@ 5.2 Moments of unbiased self-normalized importance sampling
A first result is that, for bounded test functions $f$, $\hat{F}_{u}$ has as many moments as the meeting time $\tau$, which is nearly $p$ under Assumption 2. The proof is in Appendix A.5.1.
@@
Assume that $|f|_{\infty}\leq 1$ and let $s\geq 1$. If Assumptions 1-2 hold with $p\geq 2$ and $p>s$, then the meeting time $\tau$ has $s$ finite moments, and the unbiased self-normalized importance sampling (UIS) estimator $\hat{F}_{u}$ in (36) has $s$ finite moments.
@@
Two finite moments ($s=2$ in the above results) are particularly useful. Indeed, a confidence interval for $\pi(f)$ can be constructed from independent copies of $\hat{F}_{u}$ using the Central Limit Theorem as long as $\mathbb{V}[\hat{F}_{u}]<\infty$. Beyond the empirical average, robust mean estimation strategies could be envisioned. The typical assumption in robust mean estimation is that users have access to a random variable $X$ with mean $\mu$ of interest, and a finite variance $\sigma^{2}$ that may be unknown. The goal is to aggregate independent copies of $X$ into an estimator of $\mu$ with lighter tails than the empirical average. Such results can be found in Devroye et al. (2016); Lugosi & Mendelson (2019b, a); Lecué & Lerasle (2020); Minsker & Ndaoud (2021). These techniques are directly applicable with $X=\hat{F}_{u}$ and $\mu=\pi(f)$, whereas the application of robust mean estimation to self-normalized importance sampling is not a priori straightforward, and was explored in Dau (2022) in the case of median-of-means.
@@ 5.3 The asymptotic price of bias removal
We consider the price of debiasing self-normalized importance sampling in terms of inefficiency, defined by the variance multiplied by the average cost. We start by comparing the mean squared errors of unbiased and regular IS. From (37), we take the square and use Cauchy–Schwarz to obtain
@@
The mean squared error (MSE) of IS, which is the term $\mathbb{E}[(\hat{F}(\mathbf{x}_{0})-\pi(f))^{2}]$, is of order $N^{-1}$ under conditions stated in Theorem 2.3. If we can bound $\mathbb{E}[\text{BC}^{2}]$ by a term that decreases faster than $N^{-1}$, then the MSE of UIS would be asymptotically equivalent to that of IS. Intuitively, the bias cancellation term goes to zero for two reasons: first because $\tau$ goes to one as $N\to\infty$, and the bias cancellation term equals zero in the event $\{\tau=1\}$. Secondly, each term $\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})$ goes to 00 as $N\to\infty$, under the conditions of Proposition 5.2. We obtain the following result, proven in Appendix A.5.4.
@@
The cost of UIS is that of running Algorithm 5. It starts with two draws from $\bar{q}$, i.e. $2N$ draws from $q$, and as many evaluations of the weight function $\omega$. Then either $\tau=1$ or the algorithm enters its while loop up to the meeting time $\tau$, drawing $N$ new particles at each iterate of the loop. Counting the cost in terms of the number of evaluations of $\pi$, UIS has an overall cost of $\mathcal{C}=2N+N(\tau-1)$. If Assumption 2 holds with $p\geq 2$, using Proposition 4.3 then $\mathbb{E}[\tau]=\sum_{t\geq 0}\mathbb{P}(\tau>t)$ is finite, and furthermore, $\mathbb{E}[\mathcal{C}]\sim_{N\to\infty}2N$. We summarise these first observations in the next statement.
@@
The cost of the UIS estimator $\hat{F}_{u}$ in (36) is $\mathcal{C}=2N+N(\tau-1)$, and, under Assumption 2 for $p\geq 2$, there exists a constant $C$ such that $\mathbb{E}[\tau]\leq 1+C/\sqrt{N}$.
Combining Proposition 5.4 with the above result, we see that under some conditions, the inefficiency $\mathbb{E}[\mathcal{C}]\mathbb{V}[\hat{F}_{u}]$ is equivalent to twice that of self-normalized importance sampling as $N\to\infty$. Indeed, the mean squared errors are equivalent but the cost of UIS behaves as $2N$ instead of $N$ for IS. In Algorithm 5 the $N$ particles in $\mathbf{y}_{0}$ are required to determine $\tau$ but in the event $\{\tau=1\}$ they do not participate directly in the estimator $\hat{F}_{u}$.
@@ 5.4 An improved unbiased estimator
A trick provides a remedy, and cuts the asymptotic inefficiency by a half. We can view $\hat{F}_{u}$ as a deterministic function of initial states $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ drawn from $\bar{q}$ independently, as well as extra variables (a sequence of proposals from $\bar{q}$, a sequence of uniform random variables) that we collectively label $\zeta$; we denote $\hat{F}_{u}$ by $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$. Then we define the symmetrized UIS estimator:
$\tilde{F}_{u}=\frac{1}{2}\left(\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)+\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)\right).$ (40)
Computing (40) only requires simple modifications of Algorithm 5. Indeed, either $\hat{Z}(\mathbf{x}_{0})\geq\hat{Z}(\mathbf{y}_{0})$ or $\hat{Z}(\mathbf{x}_{0})<\hat{Z}(\mathbf{y}_{0})$. In the first case, we always have $\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)=\hat{F}(\mathbf{y}_{0})$, and $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$ can be computed following Algorithm 5 and (36). In the second case, we always have $\hat{F}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)=\hat{F}(\mathbf{x}_{0})$, and $\hat{F}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)$ can be computed following Algorithm 5 and (36) with the role of $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ swapped. That trick amounts to a Rao–Blackwellization over the arbitrary specification of which draws from $\bar{q}$ are used as $\mathbf{x}_{0}$ and as $\mathbf{y}_{0}$. The following statement is a mild variation of the previous results and is stated without a proof.

### v2

@@ 4.2 Polynomial convergence rates
Furthermore, still under Assumption 2, there exists a constant $C$, independent of $t$ and $N$, such that for all $N\geq 1$ and $t\geq 0$,
@@
where the first inequality is Cauchy–Schwarz and the second uses Proposition 1.1 under Assumption 2. Furthermore, in the large $N$ asymptotics we expect $1-\hat{Z}(\mathbf{x})$ to behave as a Normal distribution with mean zero and standard deviation $\sqrt{q(\omega^{2})-1}/\sqrt{N}$, so that the expectation of its absolute value should indeed behave as $\sqrt{(2/\pi)(q(\omega^{2})-1)}/\sqrt{N}$.
@@
$\left|P^{t}(\mathbf{x},\cdot)-\bar{\pi}\right|_{\text{TV}}\leq(r(\mathbf{x}))^{t}+\frac{\tilde{C}}{\sqrt{N}\;t^{p-1}}.$ (30)
@@
$\left|P^{t}(x,\cdot)-\pi\right|_{\text{TV}}\leq(r(x))^{t}+\frac{D}{t^{p-1}}.$ (31)
Remark 4.3.
The weight $\omega$ can be unbounded while having infinitely many moments under $q$, i.e. $q(\omega^{p})<\infty$ for all $p\geq 1$. For example, this happens when $\pi$ is Gamma$(2,1)$ and $q$ is Exponential$(1)$, leading to $\omega(x)=x$. In that case Mengersen & Tweedie (1996) prove that the IMH chain cannot be geometrically ergodic, while Proposition 4.4 holds with any $p>1$. Indeed the actual decay of $|P^{t}(x,\cdot)-\pi|_{\textsc{TV}}$ could be between geometric and polynomial in $t$, for example of the form $\exp(-t^{1/2})$.
The purpose of the following example is to demonstrate that the rate $t^{-(p-1)}$ in Corollary 4.1 and Proposition 4.4 cannot be improved beyond polylogarithmic factors, without further assumptions. The proof is provided in Appendix A.4.4.
@@ 4.4 Comparing the biases of IS and IMH
In response to the first question in Section 1.1, we interpret the results on the bias of IMH and IS as follows.
  • One approach to the first question is sampling-importance resampling (SIR), which refers to the following procedure. First, draw $x_{1},\ldots,x_{N}$ independently from $q$. Compute the normalized weights $\tilde{\omega}(x_{n})=\omega(x_{n})/\sum_{m=1}^{N}\omega(x_{m})$ for $n=1,\ldots,N$. Then, draw $k\sim\text{Categorical}(\tilde{\omega}(x_{1}),\ldots,\tilde{\omega}(x_{N}))$ and return $x_{k}$. For a test function $f$ with $|f|_{\infty}\leq 1$, under the conditions of Theorem 2.1 the marginal distribution $\mu_{N}^{\text{SIR}}$ of $x_{k}$ satifies
    $\mu_{N}^{\text{SIR}}(f)-\pi(f)\sim_{N\to\infty}-q(\omega^{2}\cdot(f-\pi(f)))N^{-1}$ (34)
    Skare et al. (2003) obtains a similar result by considering the difference between the probability density function of the selected sample $x_{k}$ relative to the target $\pi$, but their approach requires stronger conditions on the weight function. They also propose a simple modification that results in a smaller bias in $N^{-2}$.
  • On the other hand, Proposition 4.4 suggests that IMH, with one proposal per iteration, after $N$ iterations, provides a sample from a distribution $qP^{N-1}$, for which, under the condition $q(\omega^{p})<\infty$ with $p>1$, for a finite constant $C$,
@@
Thus, the terminal sample after $N$ iterations of IMH is closer to $\pi$ in total variation than the sample obtained from SIR, as soon as Assumption 2 holds with $p>2$ and as $N\to\infty$. Compared to the modified SIR of Skare et al. (2003), IMH is still preferable as soon as Assumption 2 holds with $p>3$. In the case of bounded weights, MCMC methods such as PIMH or particle Gibbs (Andrieu et al., 2010) are geometrically ergodic (e.g. Lee et al., 2020) and the bias comparison is clearly at the advantage of MCMC algorithms, as discussed in Cardoso et al. (2022). This is in contrast to the comparison of asymptotic variances, which is at the advantage of IS as recalled in (6).
@@ 5.1 Construction
The bias of importance sampling was described in Section 2, and that of IMH in Section 4. Here we consider the removal of the bias, and the associated cost. There are multiple reasons to pursue bias removal for importance sampling. For example, gradient estimators in stochastic optimization, or estimators in the expectation step of the Expectation-Maximization (EM) algorithm, may be obtained by self-normalized importance sampling (e.g. Naesseth et al., 2020; Dhaka et al., 2021; Batardière et al., 2025). The resulting bias typically complicates the analysis of the convergence of encompassing optimization procedure. Another motivation for bias removal stems from the robust mean estimation literature, as described in Section 5.5. For bias removal, Middleton et al. (2019) employ common random numbers couplings of PIMH and the approach of Glynn & Rhee (2014). We pursue this strategy as well.
Upon running Algorithm 5 with $N\geq 1$, with $\tau=\inf\{t\geq 1:\mathbf{x}_{t}=\mathbf{y}_{t-1}\}$, one can compute the following unbiased estimator, denoted by UIS:
@@
(36) is an instance of unbiased MCMC (Jacob et al., 2020; Atchadé & Jacob, 2024), and various generic improvements could be considered, such as increasing the lag between the chains, or introducing a burn-in parameter. However, in the particular case of PIMH, the number of particles $N$ is a key parameter and here we focus on the regime $N\to\infty$, in which case $\hat{F}_{u}$ naturally compares with $\hat{F}(\mathbf{x}_{0})$, the regular IS estimator. Hence we view (36) as unbiased self-normalized importance sampling (UIS).
@@ 5.2 Moments of unbiased self-normalized importance sampling
A first result is that, for bounded test functions $f$, $\hat{F}_{u}$ has as many moments as the meeting time $\tau$, which is up to $p$ (non-included) under Assumption 2. The proof is in Appendix A.5.1.
@@
Assume that $|f|_{\infty}\leq 1$ and let $s\geq 1$. If Assumptions 1-2 hold with $p\geq 2$ and $p>s$, then the meeting time $\tau$ has $s$ finite moments, and the unbiased self-normalized importance sampling (UIS) estimator $\hat{F}_{u}$ in (36) has $s$ finite moments for any $N\geq 1$.
@@
Remark 5.2.
In relation to the second question in Section 1.1, the construction of UIS in (36) can be generalized to the case where $(\omega_{n},f_{n})$ jointly follow a distribution $p$ on $\mathbb{R}_{+}\times\mathbb{R}$, rather than being deterministic functions $\omega$ and $f$ of a common random variable $x\sim q$. This answers the question positively, under moment conditions on $p$.
These general results on moment bounds, particularly the case of finite second moments ($s=2$), motivate the exploration of robust estimation strategies detailed in Section 5.5.
@@ 5.3 The asymptotic price of bias removal
We consider the price of debiasing self-normalized importance sampling in terms of inefficiency, here defined by the mean squared error multiplied by the average cost (see e.g. Glynn & Whitt, 1992). We start by comparing the mean squared errors of unbiased and regular IS. From (37), we take the square and use Cauchy–Schwarz to obtain
@@
The mean squared error (MSE) of IS, which is the term $\mathbb{E}[(\hat{F}(\mathbf{x}_{0})-\pi(f))^{2}]$, is of order $N^{-1}$ under conditions stated in Theorem 2.3. If we can bound $\mathbb{E}[\text{BC}^{2}]$ by a term that decreases faster than $N^{-1}$, then the MSE of UIS would be asymptotically equivalent to that of IS. Intuitively, the bias cancellation term BC in (38) goes to zero for two reasons: first because $\tau$ goes to one as $N\to\infty$, and the bias cancellation term equals zero in the event $\{\tau=1\}$. Secondly, each term $\Delta_{t}=\hat{F}(\mathbf{x}_{t})-\hat{F}(\mathbf{y}_{t-1})$ goes to 00 as $N\to\infty$, under the conditions of Proposition 5.2. We obtain the following result, proven in Appendix A.5.4.
@@
The cost of UIS is that of running Algorithm 5. It starts with two draws from $\bar{q}$, i.e. $2N$ draws from $q$, and as many evaluations of the weight function $\omega$. Then either $\tau=1$ or the algorithm enters its while loop up to the meeting time $\tau$, drawing $N$ new particles at each iterate of the loop. Counting the cost in units of number of evaluations of $\pi$, UIS has an overall cost of $\mathcal{C}=2N+N(\tau-1)$. If Assumption 2 holds with $p\geq 2$, using Proposition 4.3 then $\mathbb{E}[\tau-1]\leq CN^{-1/2}$ for a finite constant $C$. Thus, as $N\to\infty$, $\mathbb{E}[\mathcal{C}]$ is equivalent to $2N$. We can then compare the asymptotic inefficiencies of UIS and IS as in the next statement.
@@
Denote the cost of the UIS estimator $\hat{F}_{u}$ in (36) by $\mathcal{C}=2N+N(\tau-1)$. Under the conditions of Proposition 5.4, the inefficiency (expected cost multiplied by mean squared error) of UIS is twice that of IS as $N\to\infty$:
$\lim_{N\to\infty}\mathbb{E}[\mathcal{C}]\cdot\mathbb{E}\left[(\hat{F}_{u}-\pi(f))^{2}\right]=2\times\lim_{N\to\infty}N\cdot\mathbb{E}_{\mathbf{x}\sim\bar{q}}\left[(\hat{F}(\mathbf{x})-\pi(f))^{2}\right].$
The reason for the efficiency loss in UIS is that, in Algorithm 5 the $N$ particles in $\mathbf{y}_{0}$ are required to determine $\tau$ but in the event $\{\tau=1\}$, which is increasingly likely as $N\to\infty$, these $N$ particles do not participate directly in the estimator $\hat{F}_{u}$.
@@ 5.4 An improved unbiased estimator
A simple trick provides a remedy, and cuts the asymptotic inefficiency by a half. We can view $\hat{F}_{u}$ as a deterministic function of initial states $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ drawn from $\bar{q}$ independently, as well as additional variables: an arbitrary long sequence of proposals from $\bar{q}$, and a sequence of uniform random variables used to accept or reject proposals, of the same length. Denote these two sequences by $\zeta$. Then $\hat{F}_{u}$ can be written as $\mathcal{A}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$, where $\mathcal{A}_{u}$ is now a deterministic function. Then we define the Symmetrized UIS (SUIS) estimator:
$\tilde{F}_{u}=\frac{1}{2}\left(\mathcal{A}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)+\mathcal{A}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)\right).$ (40)
Computing (40) only requires simple modifications of Algorithm 5. Indeed, either $\hat{Z}(\mathbf{x}_{0})\geq\hat{Z}(\mathbf{y}_{0})$ or $\hat{Z}(\mathbf{x}_{0})<\hat{Z}(\mathbf{y}_{0})$. In the first case, we always have $\mathcal{A}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)=\hat{F}(\mathbf{y}_{0})$, and $\mathcal{A}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)$ can be computed following Algorithm 5 and (36). In the second case, we always have $\mathcal{A}_{u}(\mathbf{x}_{0},\mathbf{y}_{0},\zeta)=\hat{F}(\mathbf{x}_{0})$, and $\mathcal{A}_{u}(\mathbf{y}_{0},\mathbf{x}_{0},\zeta)$ can be computed following Algorithm 5 and (36) with the role of $\mathbf{x}_{0}$ and $\mathbf{y}_{0}$ swapped. That trick amounts to a Rao–Blackwellization over the arbitrary specification of which draws from $\bar{q}$ are used as $\mathbf{x}_{0}$ or as $\mathbf{y}_{0}$. The following statement is a mild variation of the previous results and is stated without a proof.